At 12:14am UK local time, our monitoring reported an outage on two Vodafone provided fibre circuits delivered into our BOL8 location. These are both delivered over the same fibre bundle. These circuits provide our main office telephone service, and one of our metro ethernet fibre ring legs. After internal investigation… Continue reading →
We have just received an emergency maintenance notification from VirginMedia regarding one of our fibre optic metro ethernet links. Virgin Media will be performing emergency maintenance between midnight tonight (25/Sep/2022) and 8am. Our core network will automatically reroute during this window so no outages should be seen, however a reduction… Continue reading →
At 12:12pm today, our monitoring reported a failure of our VirginMedia metro ethernet leg within our metro ring. We have raised this as a fault with VirginMedia for investigation. We have a resilient ring so no outage or degration of service is in effect, only a reduction in resilience.
At 10:03am this morning, we saw a collection of BGP connections drop in Equinix MA1 (Manchester). These were spread across LINX MAN1, EquinixIX, and our GTT IP Transit connection. This caused a couple of minutes of instability whilst BGP reconverged and rerouted traffic automatically. A few minutes later, we saw… Continue reading →
We have identified a failure of one of our customer facing network switches in BOL8. Any customers connected directly to DS-173 in BOL8 will be affected. Other switches and customers are unaffected. We are investigating the cause and hope to have this switch back online shortly.
We had a brief network outage at approb 2:35pm today which was caused by a misconfiguration of our Vodafone fibre circuit due to human error. This affected any traffic traversing this link (such as our LINX LON1/LON2 traffic) and some traffic between Bolton and Manchester. This should now be resolved,… Continue reading →
This morning at approx. 6am UK local time, we had a switch failure of switch DS-194. This caused an outage for any directly connected customer. It also triggered an event in another switch which could affect connectivity to some virtual machines. At approx 8:45am, this switch was recovered and all… Continue reading →
LINX (London Internet Exchange) are experiencing issues with their LON1 network today. This is causing brief periods of instability and BGP session outages across their network and has been seen across multiple switches now. As a result, we have shutdown our LON1 peering until LINX confirm the issue is resolved…. Continue reading →
At 5:13am this morning we saw a brief power event on our main UPS in BOL8 datacentre followed by Switch DS-105 going offline. Although no issues are reported surrounding the power event, any customers directly connected to switch DS-105 are currently offline. DS-105 is a legacy switch stack and not… Continue reading →
At approx 11:45pm tonight (16th Dec), one of our core switches (DS-193) crashed and rebooted. This had the same error code as the other switch earlier this week and runs the same IOS version. As a result, we will now schedule IOS upgrades on all core switches.